# MENTOR: Efficient Multimodal-Conditioned Tuning for Autoregressive Vision Generation Models

### v1

@@ Abstract
Recent text-to-image models produce high-quality results but still struggle with precise visual control, balancing multimodal inputs, and demanding extensive training for complex multimodal image generation. To address these limitations, we propose Mentor, a novel autoregressive (AR) framework for efficient Multimodal-conditionEd tuNing for auTOregRessive multimodal image generation. Mentor combines an AR image generator with a two-stage training paradigm, enabling fine-grained, token-level alignment between multimodal inputs and image outputs—without relying on auxiliary adapters or cross-attention modules. Central to our method is the two-stage training paradigm: (1) a multimodal alignment stage that establishes robust pixel and semantic-level alignment between inputs and generated tokens, followed by (2) a multimodal instruction tuning stage that balance model’s integration of multimodal inputs and enhance generation controllability. Extensive experiments demonstrate that, despite a modest model size, suboptimal base components, and limited training resources, Mentor achieves a strong balance between concept preservation and prompt following on DreamBench++ benchmark, outperforming competitive baselines. Additionally, our method also delivers superior image reconstruction fidelity, broad adaptability across multimodal tasks, and an efficient training budget compared to diffusion-based counterparts. The dataset, code, and models are available in github.com/HaozheZhao/MENTOR.
Figure 1: Versatile applications built on Mentor after simply fine-tuning on corresponding datasets.
11footnotetext: Equal contribution.
@@ 1 Introduction
Recent progress in generative models has revolutionized text-to-image(T2I) generation (Ho et al., 2020; Rombach et al., 2022b; Podell et al., 2023). However, real-world applications often require more than text-only prompts. To achieve high-quality image generation, e.g., fine-grained control over generated images, the models need to seamlessly integrate multi-modal inputs, such as reference images with detailed instructions. This poses significant challenges for existing diffusion models that are predominantly focused on T2I generation. To address this, researchers have recently employed Large Multimodal Models (LMMs) to encode diverse inputs into unified embeddings compatible with diffusion models (Pan et al., 2024; Sun et al., 2024c). While this approach aids in handling complex prompts for generation like interleaved image-text instructions (Sun et al., 2024c) or multimodal in-context learning (Pan et al., 2024; Ge et al., 2024), several key limitations persist when scaling to complex multimodal control:
@@
First, the randomness of diffusion processes hinders precise, deterministic control, which is essential for high-fidelity tasks, like image reconstruction(Wang et al., 2025). Second, balancing guidance from different modalities remains challenging. Existing methods frequently exhibit modality imbalance, often over-emphasizing one modality while neglecting others(Han et al., 2024). For instance, IP-Adapter (Ye et al., 2023) and Lumina-mGPT (Zhuo et al., 2024), conditioned on text and image features, may overly favor image inputs. This imbalance stems from modality gaps, architectural limitations (Zhao et al., 2023; Cao et al., 2025; Ye et al., 2023), or suboptimal training schemes (Pan et al., 2024; Han et al., 2024). Third, many diffusion-based methods (Pan et al., 2024; Sun et al., 2024c), particularly those with auxiliary alignment components (e.g., learned adapters (Pan et al., 2024), regression heads (Sun et al., 2023, 2024c), specialized embeddings (Ge et al., 2023)), demand large-scale training (Sun et al., 2024c; Pan et al., 2024; Li et al., 2023a; Ge et al., 2024), incurring substantial computational costs. These challenges raise a critical question: Is there a more efficient and controllable paradigm for balancing complex multimodal conditions in image generation?
@@
To address aforementioned limitations, we propose Mentor, a straightforward and efficient autoregressive (AR) framework for controllable multimodal image generation. Unlike diffusion models that rely on complex cross-attention layers for multimodal conditioning and demand extensive training resources (Ge et al., 2024; Sun et al., 2024c; Pan et al., 2024; Li et al., 2023a), Mentor leverages a unified transformer architecture that directly aligns multimodal inputs with output image tokens. This design inherently simplifies architecture and training, reduces the need for auxiliary components for alignment, and significantly lowers training requirements. Our framework employs a multimodal encoder to project multimodal inputs into a unified representation. An AR transformer decoder then generates image tokens deterministically, conditioned on this representation. To ensure effective and balanced modality integration, which is curial for multimodal image generation(Han et al., 2024), we adopt a two-stage training paradigm: (1) pretraining on image reconstruction and object segmentation to enable robust pixel-level and semantic alignment, followed by (2) multimodal instruction tuning with diverse tasks, such as image recovery and subject-driven generation, explicitly training the model to effectively leverage and balance varied multimodal inputs for coherent generation.
Notably, despite its simplicity and the use of suboptimal checkpoints, Mentor achieves state-of-the-art(SOTA) performance on the challenging DreamBench++ benchmark—using 10× less training data than leading baselines. It outperforms resource-intensive baselines with powerful generators like SDXL (Podell et al., 2023) and SD3 (Esser et al., 2024) by 30% while dramatically reducing computational and data demands. Controlled experiments demonstrate the advantages of Mentor over diffusion-based methods in multimodal fidelity, training efficiency, and generation controllability.
@@
Overall, our contributions are as follows: (1) A novel autoregressive framework for efficient multimodal generation, achieving, achieving SOTA performance; (2) A two-stage training strategy for multimodal-conditioned tuning, enabling robust alignment and balanced modality integration with significantly reduced training cost; (3) Extensive experiments demonstrating the superior efficiency, controllability, and fidelity of Mentor as a compelling alternative to diffusion-based methods.
@@ 2 Method
This section details an efficient autoregressive framework for controllable multimodal image generation, achieving precise image control and balancing guidance across multiple modalities with minimal cost. § 2.1 describes the autoregressive training objective for our framework. Next, § 2.2 presents a straightforward yet efficient model architecture, detailing how the framework accommodates multimodal inputs and supports autoregressive generation. Crucially, § 2.3 introduces a two-stage training paradigm aimed at balancing the influences of different modalities during image generation. Finally, § 2.4 describes our automated pipeline for scalable multimodal data curation.
2.1 Preliminary
Training Objective Our model employs teacher forcing to predict image tokens, conditioned on (i) previously generated tokens and (ii) multimodal context $\mathbf{h}$. Given the multimodal condition: $\mathbf{c}^{(0)}=\{\mathcal{I},\mathcal{T}\}$ (visual and textual inputs), a multimodal encoder $\phi$ first encodes $\mathbf{c}^{(0)}$ and subsequently uses an MLP layer to project them into space of the image decoder to form a unified representation $\mathbf{h}$:
$\mathbf{H}=\text{MLP}(\phi\bigl{(}\mathbf{c}^{(0)}\bigr{)})=(\mathbf{h}_{1},\dots,\mathbf{h}_{M})\in\mathbb{R}^{M\times d},\qquad\mathbf{h}_{j}\in\mathbb{R}^{d}.$ (1)
$\small\theta(\mathbf{y}\mid\mathbf{H})=\prod_{i=1}^{L}\theta\bigl{(}y_{i}\mid y_{<i},\mathbf{H}\bigr{)}.$ (2)
$\small\mathcal{L}_{\text{CE}}(\theta,\phi)=-\mathbb{E}_{(\mathbf{y},\mathbf{c}^{(0)})\sim\mathcal{D}}\left[\sum_{i=1}^{L}\log\theta\bigl{(}y_{i}\mid y_{<i},\mathbf{H}\bigr{)}\right].$ (3)
Classifier-free Guidance To enhance multimodal generation controllability, we apply Classifier-Free Guidance (CFG) (Sun et al., 2024b). During training, multimodal conditioning $\mathbf{H}$ is replaced by a learned unconditional embedding $\mathbf{H}_{u}$ with probability $p$. At inference time, token logits $\ell_{g}$ are recalculated by interpolating between the conditional logits $\ell_{c}$ (from $\mathbf{H}$) and unconditional logits $\ell_{u}$ (from $\mathbf{H}_{u}$), controlled by a scaling parameter $\lambda$: $\ell_{g}=\ell_{u}+(\ell_{c}-\ell_{u})\times\lambda$.
@@ 2.2 Model Design
As illustrated in Figure 3, Mentor architecture comprises two core components: a multimodal encoder and an autoregressive generation decoder. These components are designed to unify multimodal inputs into a shared embedding and generate image tokens sequentially conditioned on the unified embedding, respectively. A lightweight projection layer bridges the encoder’s output to the decoder’s input embedding space, enabling seamless integration between the two components.
Multimodal Encoder The multimodal encoder integrates multimodal inputs from frozen pretrained vision ($\phi_{V}$) and language ($\phi_{L}$) encoders into a shared latent space. This module projects visual features from $\phi_{V}$ into $\phi_{L}$’s embedding space using a lightweight connector module ($\psi$), yielding a unified multimodal representation $\mathbf{H}=(\mathbf{h}_{1},\dots,\mathbf{h}_{M})$, where $\mathbf{h}_{j}\in\mathbb{R}^{d}$. Meanwhile, a critical challenge arises from the visual tokenization, where a vision encoder produces hundreds of tokens per image. It inflates the context length and leads to substantial computational costs, especially in multi-image scenarios. Compressing visual tokens can mitigate costs, but may risk losing fine-grained visual information. To navigate this trade-off, two architectures for $\psi$ are investigated:
  • MLP-based Projection: Inspired by Liu et al. (2023), we adopt a multi-layer perceptron that directly projects visual tokens to maintain detailed visual semantics with minimal information loss.
  • Query-based Token Distillation: We use a Query-Former Li et al. (2023b) to compress lengthy visual token sequences into a fixed-size representation. To guide the distillation, we condition the Query-Former on textual queries(Li et al., 2023a; Dai et al., 2023) that highlight important concepts in images. It aims to guide the model to selectively extract semantically relevant visual features based on textual queries.
Figure 3: Overview of Mentor. Left panel illustrates the model structure, where visual and textual inputs are encoded into a unified latent to guide autoregressive image generation. Right panel highlights the two-stage training paradigm: (1) Multimodal Alignment Tuning, enabling pixel and semantic-level alignment between inputs and output tokens; and (2) Multimodal Instruction Tuning, compels model to effectively balance influence of different modalities.
Autoregressive Generation Decoder A transformer-based autoregressive decoder generates a image token sequence $\mathbf{y}=(y_{1},\dots,y_{L})$ conditioned on the prefix $\mathbf{H}$ generated by the multimodal encoder and previously generated tokens $y_{<i}$. It operates in a shared embedding space with the encoder’s output and shares the same vocabulary as the VQGAN (Esser et al., 2020) that is used for image tokenization. The generated token sequences are subsequently decoded into images using the VQGAN decoder. This unified autoregressive structure facilitates unified training via next-token prediction.
@@ 2.3 Two-Stage Training Paradigm
As highlighted in the introduction, effectively aligning disparate modalities and balancing their influence are crucial challenges for multimodal conditional generation. Our carefully designed two-stage training paradigm directly addresses these issues, moving beyond initial coarse alignment to foster robust understanding and balanced integration of diverse inputs, as illustrated in Figure 3.
Stage 1: Multimodal Alignment Tuning While the initial connector training provides preliminary semantic alignment, we observed that the model primarily interprets visual inputs semantically like text captions, neglecting crucial visual and spatial details necessary for precise image generation. This stage is dedicated to explicitly enhancing both pixel and semantic-level modality alignment and promoting the utilization of visual information. In Stage 1, we employ three complementary tasks:
  • Image reconstruction, where models must faithfully reconstruct an input image conditioned on itself, with the corresponding caption randomly provided or omitted, reinforcing pixel-level fidelity.
  • Object segmentation, where models are given an input image and a target object label and must generate an end-to-end segmented figure for that object. This task compels the model to explicitly capture fine-grained visual details and spatial structures associated with specific semantic concepts.
  • T2I generation, using image-caption pairs to preserve and reinforce foundational generative capabilities learned during pretraining of the decoder.
Importantly, incorporating the segmentation task alongside image reconstruction mitigates the risk of the model degenerating into a trivial copy-paste behavior, such as simply replicating the input. The segmentation objective compels the model to produce semantically meaningful and spatially precise visual outputs. This complementary effect has been further analyzed and validated in § 3.3.
Stage 2: Multimodal Instruction Tuning Stage 2 aims to endow the model with robust instruction-following and cross-modal reasoning capabilities for nuanced and controllable multimodal generation, building upon the alignment and visual fidelity established in Stage 1. The model is expected to jointly attend to and integrate diverse input modalities in a balanced and controllable manner. To achieve this, we employ a multimodal instruction tuning strategy based on a carefully curated mixture of training tasks. Specifically, we reuse the T2I generation and object segmentation tasks from Stage 1, maintaining identical data and formulations. These tasks respectively reinforce the model’s ability to adhere to and utilize textual and visual modalities, helping preserve foundational skills and stabilize the training process. In addition, we introduce two novel tasks specifically designed to enhance instruction adherence and foster a balanced integration of multimodal inputs, preventing the model from over-emphasizing a single modality while neglecting others:
  • Image recovery, where we synthetically distort images by rotating, resizing, and compositing segmented objects onto random backgrounds, then pair synthetic images with their original captions to create inputs. The model is then required to reconstruct the original image from the distorted input and corresponding caption. It compels the model to extract and integrate essential visual details from noisy or incomplete visual inputs while leveraging textual cues to infer and restore missing components, promoting robust multimodal reasoning and error-correction performance.
  • Subject-driven image generation, where the model is conditioned on reference image, subject label and textual instruction to generate images. It require the model to actively preserve the subject’s visual identity from the reference image while strictly adhering to the textual instructions for image generation. This task serves as a comprehensive end-to-end objective, fully exercising model’s cross-modal fusion and instruction-following abilities.
Overall, this training strategy—combining continued refinement of core capabilities with targeted instruction-based tasks—ensures the model to learn to integrate visual and textual information in a harmonious and controllable way. It mitigates over-reliance on certain modality and enables precise, controllable multimodal conditional generation, which are further discussed in § 3.3.
@@ 2.4 Data Construction
To support our two-stage training paradigm, we construct a large-scale multimodal dataset comprising approximately 3 million samples across all training tasks. It integrates open-source resources, synthetic data, and automated annotations to ensure scalability, diversity, and strong task alignment. For image reconstruction and T2I generation, we collect image-text pairs from datasets like CC12M (Changpinyo et al., 2021) and Midjourney-Niji (Emporium, 2024). To broaden domain coverage (e.g., human subjects, artistic scenes), we generate additional samples using T2I models like Flux.1 (BlackForestLabs, 2024) and Stable Diffusion v3.5 (Esser et al., 2024), with prompts generated by advanced LLMs (OpenAI, 2024) to enhance semantic and visual diversity. For segmentation and image recovery, which require fine-grained object-level annotations, we design an automated pipeline that combines state-of-the-art LMMs (Wang et al., 2024a) with segmentation models(Kirillov et al., 2023). Given an image, LMMs are queried to produce a comprehensive caption and extract a list of concrete, segmentable objects. For each object, LMMs predicts its spatial location, providing a bounding box and a set of 2D keypoints, which guide the segmentation model in producing high-quality, semantically consistent masks. For subject-driven image generation, we leverage the OminiControl dataset (Tan et al., 2025), re-captioned using LMMs to accurately extract subject-relevant descriptions. Additionally, we reverse image pairs to effectively double the usable data. Data construction and formation are detailed in Appendix D.

### v2

# MENTOR: Efficient Autoregressive Image Generation with Balanced Multimodal Control

@@ Abstract
Recent text-to-image models achieve impressive visual quality but still face challenges in precise controllability, balancing multimodal inputs, and high training cost for multimodal image generation. To address these limitations, we propose Mentor, an autoregressive (AR) framework with a two-stage training paradigm for controllable multimodal image generation: (1) a multimodal alignment stage that establishes robust pixel and semantic-level alignment between inputs and generated tokens, followed by (2) a multimodal instruction tuning stage that balances the model’s integration of multimodal inputs and enhances generation controllability. Extensive experiments on DreamBench++ and DreamBench demonstrate that, despite modest model size and training resources, Mentor achieves a strong balance between textual and visual guidance for controllable image generation, delivering competitive performance at significantly lower computational cost compared to leading baselines. Moreover, our approach attains superior image reconstruction fidelity, broad adaptability across different tasks, and training efficiency.
[…]
Figure 2: Qualitative examples of different tasks built on Mentor after simply fine-tuning.
MENTOR: Efficient Autoregressive Image Generation with Balanced Multimodal Control
Haozhe Zhao1,  Zefan Cai2,  Shuzheng Si3,  Liang Chen4,   Jiuxiang Gu,  Wen Xiao5,  Minjia Zhang1,  Junjie Hu2 1University of Illinois Urbana-Champaign 2University of Wisconsin-Madison 3Tsinghua University 4Peking University 5Microsoft haozhez6@illinois.edu haozhezhao.github.io/MENTOR.page
@@ 1 Introduction
Recent progress in generative models has revolutionized text-to-image (T2I) generation (Ho et al., 2020; Rombach et al., 2022b; Podell et al., 2023). However, real-world applications often require more than text-only prompts. To achieve controllable image generation, e.g., fine-grained control over generated images, models need to seamlessly integrate multi-modal inputs, such as a reference image together with a detailed text prompt. This poses significant challenges for existing models that are predominantly focused on T2I generation. To address this, researchers integrate Large Multimodal Models (LMMs) with generative models (Pan et al., 2024; Sun et al., 2024c; Xiao et al., 2024a; Zhuang et al., 2025) to better handle multimodal inputs. Although effective for tasks like interleaved image-text generation (Sun et al., 2024c) and multimodal in-context learning (Pan et al., 2024; Ge et al., 2024), these approaches still face three challenges when scaling to complex multimodal control, especially under limited resources:
@@
First, stochastic sampling in diffusion processes makes precise and deterministic control difficult, which is essential for high-fidelity tasks (e.g., faithful reconstruction) (Wang et al., 2025). Second, balancing guidance across modalities remains challenging. Existing methods exhibit modality imbalance, overemphasize one modality while neglecting the other (Han et al., 2024b). This phenomenon appears in both diffusion and autoregressive (AR) paradigms—for instance, IP-Adapter (Ye et al., 2023) and Lumina-mGPT (Zhuo et al., 2024), conditioned on text and image features, tend to favor image guidance. Such imbalance may stem from modality gaps, architectural limitations (Zhao et al., 2023; Cao et al., 2025; Ye et al., 2023), or suboptimal training schemes (Pan et al., 2024; Han et al., 2024b). Third, existing methods rely on auxiliary alignment components, such as learned adapters (Pan et al., 2024), regression heads (Sun et al., 2023, 2024c) or specialized embeddings (Ge et al., 2023)), and demand large-scale training (Sun et al., 2024c; Pan et al., 2024; Ge et al., 2024), leading to significant computational costs. These observations motivate a question: Is it possible to design an efficient framework for controllable multimodal image generation under limited resources?
@@
To address these limitations, we propose Mentor, an efficient autoregressive (AR) framework for controllable multimodal image generation. Unlike diffusion models that demand complex cross-attention layers for multimodal conditioning and extensive training resources (Sun et al., 2024c; Pan et al., 2024; Li et al., 2023a), Mentor employs a unified transformer that directly aligns multimodal inputs with output tokens, simplifying the architecture, removing auxiliary alignment modules, and substantially reducing training costs. Our framework uses a multimodal encoder to project inputs into a unified representation, which a transformer decoder then uses to deterministically generate image tokens. To ensure effective and balanced modality integration (Han et al., 2024b), we further adopt a two-stage training paradigm: (1) a multimodal alignment stage that builds robust pixel- and semantic-level alignment between inputs and generated tokens, followed by (2) a multimodal instruction tuning stage that balances the modality fusion and enhances generation controllability.
Notably, despite its simplicity and usage of suboptimal checkpoints, Mentor achieves competitive performance on various benchmarks including Dreambench (Ruiz et al., 2023) and Dreambench++ (Peng et al., 2025) with over 10× less resources than leading baselines. It surpasses resource-intensive models equipped with powerful generators such as SDXL (Podell et al., 2023), SD3 (Esser et al., 2024) and Infinity (Han et al., 2024a). Controlled experiments show that Mentor offers a favorable trade-off between efficiency and multimodal balance. Further analyses highlight its adaptability across reinforcement learning and diverse multimodal tasks, offering a practical framework for controllable multimodal generation.
@@
Overall, our contributions are as follows: (1) An autoregressive framework for efficient controllable multimodal image generation; (2) A two-stage training strategy that enables robust alignment and balanced modality integration with substantially reduced computational cost; (3) Experiments demonstrating the superior efficiency, controllability, and fidelity of Mentor as a compelling framework for controllable multimodal generation.
Figure 3: Overview of Mentor. Left panel illustrates model structure, where visual and textual inputs are encoded into a unified latent to guide autoregressive image generation. Right panel highlights two-stage training paradigm: (1) Multimodal Alignment Tuning, enabling pixel and semantic-level alignment between inputs and output tokens; and (2) Multimodal Instruction Tuning, compels model to effectively balance influence of different modalities.
@@ 2.1 Model Design
As illustrated in Figure˜3, Mentor architecture comprises two core components: a multimodal encoder and an autoregressive generation decoder. These components are designed to unify multimodal inputs into a shared embedding and generate image tokens sequentially conditioned on the unified embedding, respectively.
Multimodal Encoder The multimodal encoder integrates multimodal inputs from frozen pretrained vision ($\phi_{V}$) and language ($\phi_{L}$) encoders into a shared latent space. This module projects visual features from $\phi_{V}$ into $\phi_{L}$’s embedding space using a lightweight connector module ($\psi$), yielding a unified multimodal representation $\mathbf{H}=(\mathbf{h}_{1},\dots,\mathbf{h}_{M})$, where $\mathbf{h}_{j}\in\mathbb{R}^{d}$. In practice, we adopt a multi-layer perceptron layer as the connector module that directly projects visual tokens to maintain detail information.
Autoregressive Decoder A transformer-based autoregressive decoder generates a image token sequence $\mathbf{y}=(y_{1},\dots,y_{L})$ conditioned on the prefix $\mathbf{H}$ generated by the multimodal encoder and previously generated tokens $y_{<i}$. It operates in a shared embedding space with the encoder’s output and shares the same vocabulary as the VQGAN (Esser et al., 2020) that is used for image tokenization. The generated token sequences are subsequently decoded into images using the VQGAN decoder. This unified autoregressive structure facilitates unified training via next-token prediction.
@@ 2.2 Two-Stage Training Paradigm
Effectively aligning different modalities and balancing their influence are crucial challenges for multimodal image generation. Our two-stage training paradigm directly addresses these issues, moving beyond initial coarse alignment to foster robust understanding and balanced integration of diverse inputs, as illustrated in Figure˜3.
Stage 1: Multimodal Alignment Tuning. Although the multimodal encoder can encode multimodal inputs for the generator, we observe that the model tends to interpret visual information semantically like text captions, while neglecting visual details. To explicitly strengthen both pixel- and semantic-level alignment and promote richer visual utilization, we introduce three complementary tasks: (1) image reconstruction, where the model reconstructs an input image (with or without corresponding caption) to enhance pixel-level fidelity; (2) object segmentation, where the model generates an object-specific segmentation mask given an image and target label, enforcing attention to fine-grained visual details and spatial structures associated with semantic concepts; and (3) text-to-image (T2I) generation, using image–caption pairs to preserve and reinforce generative capability. Notably, combining segmentation with reconstruction prevents trivial copy behavior (i.e., simply replicating input images) by encouraging semantically meaningful and spatially precise outputs. This complementary effect is further analyzed in §˜3.3.
Stage 2: Multimodal Instruction Tuning. Building upon the alignment and visual fidelity established in Stage 1, this stage aims to equip the model with robust instruction following and cross-modal reasoning abilities. To encourage the model to integrate different modalities in a balanced manner, we adopt a multimodal instruction tuning strategy using a curated mixture of tasks. We retain the text-to-image (T2I) generation and object segmentation tasks from Stage 1, preserving their formulations to maintain foundational skills and stabilize training. In addition, two new tasks are introduced to enhance instruction adherence and balanced modality integration. (1) image recovery introduces synthetic distortions—such as rotation, resizing, and compositing segmented objects onto random backgrounds—paired with original captions. The model are required to reconstruct the original image from the distorted input and caption, encouraging it to extract essential visual cues and leverage textual information to restore missing details. (2) subject-driven image generation conditions the model on a reference image, subject label, and textual instruction to generate new images. This task requires preserving the subject’s visual identity while faithfully following textual directives, serving as a comprehensive end-to-end objective. Overall, this stage enables the model to integrate visual and textual information in a balanced manner, effectively mitigating over-reliance on a single modality and supporting precise multimodal generation, as further discussed in Table˜3.
@@ 2.3 Data Construction
To support our two-stage training paradigm, we construct a multimodal dataset of approximately 3 million samples. The dataset integrates open-source resources, synthetic data, and automated annotations to ensure scalability and diversity. For image reconstruction and T2I generation, we collect image–text pairs from datasets such as CC12M (Changpinyo et al., 2021) and Midjourney-Niji (Emporium, 2024). To expand domain coverage (e.g., human subjects, artistic scenes), we further synthesize samples using T2I models like Flux.1 (BlackForestLabs, 2024), with prompts generated by advanced LLMs (OpenAI, 2024) to enhance semantic and visual diversity. For segmentation and image recovery, which require fine-grained object-level annotations, we design an automated pipeline that combines state-of-the-art LMMs (Wang et al., 2024a) with segmentation models(Kirillov et al., 2023). For subject-driven image generation, we leverage the OminiControl dataset (Tan et al., 2025), re-captioned using LMMs to accurately extract subject-relevant descriptions. Additionally, we reverse image pairs to effectively double the usable data. Data pipeline and formation are detailed in Appendix˜D.